OpenStudy (anonymous):

72 is what percent of 480

OpenStudy (anonymous):

15%

OpenStudy (anonymous):

72/480=.15

OpenStudy (mathstudent55):

Translate the words into an equation. Use these rules: "is" means = "of" means * (times) "percent" means /100 (divided by 100) "what" means x (the variable x) 72 is what percent of 480 72 = x/100 * 480 \(72 = \dfrac{x}{100} \times 480 \) Now solve the equation for x.